The temperature of a 12-g sample of lead metal increases from 22 C to 35 C upon the addition of 20.0 J of heat Determine the specific heat capacity of lead. (5 points)
Answer #1

ans)

from above data that

we know that

applyheat absdorbed Q = mCDT

given that temperatures=35-22

m=12 g

q=20 j

here we need C

so C = Q/mDt

now bys ubsuitituning all values

C = 20/(12* (35-22)

C = 0.128 approximately=0.13